OpenStudy (anonymous):

use Gauss's approach to find the follwoing sums (do not use formulas). a... 1+2+3+4+...+999 b... 1+3+5+7+...+997 the sum of sequance a is... the sum of sequance b is...

OpenStudy (anonymous):

guys... I'm in precal, I don't know how to do this stuff yet

OpenStudy (anonymous):

What I find amusing about this, is that Gauss's approach actually was a formula, so I wonder why they mention not to use a formula.

OpenStudy (anonymous):

gauss approach is to add forwards and backwards

OpenStudy (anonymous):

\[S=1+2+3+4+...+997+998+999\]\[S=999+998+997+...+3+2+1\] \[2S=1000+1000+1000+...+1000=999\times 1000\] \[S=\frac{999\times 1000}{2}\] at least that is the legend